OpenStudy (anonymous):

What is the simples form for 88%

OpenStudy (anonymous):

88% as a fraction is: \[\frac{ 88 }{ 100 }\] Then divide by both of their similar GCF's (Greatest common factor) to get the simplest form: \[\frac{ 22 }{ 25 }\]

OpenStudy (anonymous):

\[\frac{ 88 }{ 100 }or0.88\]